OpenStudy (anonymous):

Find the coordinates of the point where the folium has a vertical tangent. Folium of Descartes, x^3+y^3-9xy=0

OpenStudy (anonymous):

\[3x^2 + 3y^2 \frac{dy}{dx} - 9(x \frac{dy}{dx} + y) =0\]

OpenStudy (anonymous):

\[\frac{dy}{dx}(3y^2 -9x) = (9y-3x^2) \]

OpenStudy (anonymous):

\[\frac{dy}{dx}= \frac{ 9y-3x^2}{3y^2-9x}\]

OpenStudy (anonymous):

This derivative is undefined when denominator equals zero.

OpenStudy (anonymous):

\[(y \sqrt{3} - 3\sqrt{x})(y \sqrt{3}+3\sqrt{x}) =0\]

OpenStudy (anonymous):

\[y= +- \sqrt{3x}\]

OpenStudy (anonymous):

In terms of substituting that condition into the given equation above. mmmm that will be difficult to find an exact solution for x.